# Supermassive black hole growth from stellar binary encounters

### v2

@@ Abstract
The growth of supermassive black holes (SMBHs) remains a central problem in astrophysics, with current observations providing only limited constraints on the underlying mechanisms. One possible growth channel is stellar accretion via the Hill’s mechanism, wherein a SMBH tidally breaks up a passing binary star, capturing and eventually accreting a member of the binary. We adopt a model to predict capture rates from parameters that include the central number density of stars, the stellar velocity dispersion, the binary fraction, and black hole mass. We then use this model to estimate the growth of SMBHs across a range of galactic environments. In a data set of over 90 galaxies of various types and masses, we identify two candidates with SMBHs for which stellar accretion may be a driver of recent growth. Closer to home, a recent analysis of observed hypervelocity stars from the Large Magellanic Cloud (LMC) implicates binary star interactions with a massive black hole. Every hypervelocity star produced in this way leaves a bound partner that may be accreted, providing an active growth channel for the LMC’s black hole.
@@ 1 Introduction
Enshrouding millions to billions of solar masses within an event horizon of remarkably small size, supermassive black holes (SMBHs) are famously elusive. Yet observations have teased out irrefutable evidence for their existence. In our own galaxy, years-long tracking of stars on tight orbits about a common point beautifully reveal Sgr A, the Milky Way’s four million solar-mass black hole (Ghez et al., 2008; Genzel et al., 2010). The Event Horizon Telescope has since mapped gas flows near the supermassive black hole, unveiling a “shadow” surrounding the event horizon (Event Horizon Telescope Collaboration et al., 2022).
@@
Before these achievements were technologically feasible, Hills (1988) sought other, circumstantial evidence for supermassive black holes. Focusing on the Milky Way, Hills calculated that a binary star passing close to a SMBH in the center of the Galaxy woud be broken up, with one stellar component flung out at high speed in a gravitational sling-shot, while its partner would remain bound to the black hole. The unbound star would be traveling fast enough to escape the Galaxy. Detection of a hypervelocity star (HVS) produced this binary disruption (“break-up”) mechanism would be compelling evidence of the black hole itself.
@@
The discovery of a hypervelocity star by Brown et al. (2005) confirmed Hills’ prediction that a supermassive black hole in the center of the Milky Way produces fast stars with no other plausible origin. Subsequent detections (e.g., Brown et al., 2006, 2009; Zheng et al., 2014, see also Verberne et al. 2025) built up a census of HVSs. The HVS population thus contributed to the growing body of evidence demonstrating that Sgr A was indeed a SMBH. More recently, and based on astrometry from Gaia (Gaia Collaboration et al., 2023), some of these same stars now appear to originate from the Large Magellanic Cloud (LMC). Han et al. (2025) analyzed these sources to infer the existence of LMC, a $\sim{6}\times 10^{5}$ $\rm M_{\odot}$ black hole in the dwarf galaxy. The Hills mechanism is fulfilling its promise as a signpost for SMBHs.
@@
Hypervelocity stars are the lead astronomical actors in the Hills mechanism, yet they are only half the story. The bound partner in each binary break-up remains close to the black hole, and eventually may be accreted onto it (e.g., Bromley et al., 2012). A signature of this process is a tidal disruption event (TDE), which occurs when a black hole’s tidal forces shred the star (Hills, 1975; Frank & Rees, 1976; Lightman & Shapiro, 1977; Rees, 1988, see also Gezari 2021 and references therein). This outcome is possible for black holes with masses less than $\sim{10}^{8}$ $\rm M_{\odot}$, for which the typical tidal disruption radius is outside the event horizon. With more massive black holes, the star simply strays across the event horizon without a trace.
@@
Any bound star can be consumed, either tidally disrupted or swallowed by a black hole, once its orbital angular momentum drifts below a critical value, allowing it to plunge toward the event horizon (e.g., Merritt, 2013; Chen et al., 2023). As fuel for SMBHs, higher-mass stars are more susceptible to tidal disruption because of their diffuse outer layers, which experience weaker gravitational binding forces. As a result, when these stars approach an SMBH, their outer envelopes are easily stripped away by tidal forces (Chen et al., 2023). These partial disruptions can occur multiple times before a massive star is completely destroyed (Ryu et al., 2022). In any case, these stars, as well as their lower-mass counterparts, may ultimately be accreted by the SMBH (e.g., Thomsen et al., 2022; Webb & Foustoul, 2022; Olejak et al., 2025). Stellar accretion may indeed be a dominant early-stage growth channel for black holes in the mass range of $10^{2}$$10^{6}$ $\rm M_{\odot}$, including single-star captures (Stone et al., 2017) and binary break-ups (González Prieto et al., 2025).
@@
A variety of mechanisms have been proposed to explain the growth of supermassive black holes, including mergers and gas accretion (e.g., Silk & Rees, 1998; Yu & Tremaine, 2002; Di Matteo et al., 2003; Montesinos Armijo & de Freitas Pacheco, 2011; Ryu et al., 2022). Yet no single channel has emerged to account for the evolution of all SMBHs. Here, our focus lies on the role of binary stars and the Hills mechanism as contributors to SMBH growth. We begin in §2 by revisiting the theoretical description of this growth channel from Bromley et al. (2012), implementing it in a Python code to track binary star interaction rates with a SMBH and the subsequent evolution of the black hole mass. In §3, we apply the model to galactic data provided by Hannah et al. (2024), which span a wide range of black hole masses and galactic host properties. One feature of this analysis is a prediction for the relationship between black hole mass and central stellar densities, both of which are critical to the rate of disruption of stellar binaries. Closer to home, we assess the growth of LMC, the black hole candidate in the Large Magellanic Cloud S4. We summarize our results in §5.
@@ 2 The model: black hole growth through the Hills mechanism
At larger closest-approach distances, tidal forces from the black hole are weaker. The gravitational influence of the binary at $r_{\rm close}$ is characterized by
$r_{\rm D}\equiv{r_{\rm close}}\left[\frac{m_{\rm bin}}{11M_{\bullet}}\right]^{1/3}$ (1)
where $m_{\rm bin}$ is the total binary mass (Hills, 1988). If this measure is greater than the binary’s orbital separation $a_{\rm bin}$, then the binary generally will survive its encounter with the SMBH. [^1] The quantity $r_{\rm D}$ is related to the Hill radius, $r_{\rm H}\approx 1.54r_{\rm D}$, from planetary dynamics, describing the gravitational influence of a smaller body on a circular orbit around a larger one. Our choice is also similar to J. Hills’ parameter $D$, which is approximately $175r_{\rm D}/a_{\rm bin}$. For example, a stellar binary with total mass of 2 $\rm M_{\odot}$and a semimajor axis of 1 au can safely pass a $10^{6}$ $\rm M_{\odot}$ black hole if $r_{\rm close}$ is greater than 175 au.
At intermediate closest approach distances, between stellar disruption and $r_{\rm D}$, outcomes depend on details of the orbits including binary separation, orientation, and orbital phase, as well the masses of the two stars and the black hole. A key to Hills’ approach is that these details are treated probabilistically. On the basis of numerical trials with randomized orbital configurations, Hills quantified the probability that a binary encounter with a SMBH will lead to binary break-up entirely in terms of $r_{\rm D}$ and the binary’s orbital separation:
$P_{\text{bbu}}\approx 1-\frac{r_{\rm D}}{a_{\rm bin}}.$ (2)
Distributions of binary properties in a stellar population can inform the range of break-up outcomes, and these outcomes can further be translated into likely HVS ejection speeds and captured star orbital elements based on energy and angular momentum conservation (Hills, 1988; Bromley et al., 2006).
@@ 2.1 Binary break-up rates
Here we are interested in average or representative binary properties to predict the change in mass of a SMBH from the accretion of stars captured by the Hills mechanism. Our starting point for the rate is $nv\sigma$” from kinematics, where $n$ is the number density of binaries, $v$ is their mean speed in a region surrounding a SMBH, and $\sigma$ is the cross-sectional area of the black hole — the size of the hypothetical target for a binary star to hit in order to be split into a hypervelocity star and a bound partner. We consider each factor in turn.
@@
where $m_{\rm avg}\approx$ 0.3 $\rm M_{\odot}$ is the average mass of a star, counting binary components as individual stars. In the Milky Way, the binary fraction overall is roughly half of all stars (Duchêne & Kraus, 2013, and references therein), though it is dependent on spectral type of the components (e.g., Moe & Di Stefano, 2017). Typically younger, more massive stars have a higher binary fraction than less massive, older stars.
The central density of stars, $\rho_{0}$, depends strongly on the SMBH mass, $M_{\bullet}$. For large elliptical galaxies, the density falls with increasing black hole mass $\rho_{0}\sim 1/M_{\bullet}$ (e.g., Faber et al., 1997). The trend is opposite in smaller galaxies. Both central stellar density and black hole mass increase with the size and mass of the galactic host. Here, we introduce the parameterization,
@@
where $\rho_{0}$ is the central density within some small (e.g., 5 pc) of the SMBH, and $_{b}\approx 4\times 10^{7} M\rm M_{\odot}$  is the black hole mass associated with the break between two power-law relationships; $\rho_{0}\sim M_{\bullet}$ in the low-mass regime and $\rho_{0}\sim 1/M_{\bullet}$ for masses much greater than $M_{b}$. In the next section, We demonstrate the effectiveness of Equation (4) using data kindly provided by Hannah et al. (2024) that inspired our choice of parameterization.
The central velocity dispersion of stars in the cores of galaxies, $v_{\rm disp}$, is typically of the order of 100 km/s, with values that increase with SMBH mass and total galactic mass (e.g., Faber et al., 1997; Bernardi et al., 2003; McConnell & Ma, 2013; Greene et al., 2020). The range of observed values is narrow compared with that of the masses of known SMBHs, because $v_{\rm disp}$ trends as the fourth root of $M_{\bullet}$ (Kormendy & Ho, 2013) and broader galactic properties including luminosity (e.g., Faber & Jackson, 1976), both of which can span many orders of magnitude.
The third factor in the kinematic description is $\sigma_{\bullet}$, the cross-sectional area of the black hole in scattering events with binaries. This term depends on uncertain quantities including the distribution of binary orbital distances/semimajor exes and the nature of their orbits around the black hole. Following our review of the Hill mechanism above, we adopt $r_{\rm close}$ corresponding to values where the break-up $P_{\text{bbu}}\gtrsim 0.5$ as a measure of the cross-section near the black hole.
The strongest impact of black hole mass is the boost to its cross-sectional area by the long-range effect of gravity on approaching stars. The aim of a binary toward the black hole may be poor, yet gravity draws it in. The cross section is increased by a gravitational focusing factor
@@
(Bondi & Hoyle, 1944).
Putting these factors together to derive an $nv\sigma$” interaction rate, we find that the break-up rate of binary stars through the Hills mechanism is
$k$ $=$ $\frac{\rho_{0}}{m_{\rm avg}}\frac{f_{\rm bin}}{(1+f_{\rm bin})}\cdot v_{\rm disp}\cdot\pi f_{\text{grav}}r_{\rm close}^{2}$ (6)
$\approx$ $0.0011\left[\frac{\rho_{0}}{10^{4}\ \text{\footnotesize$\rm M_{\odot}$/pc}^{3}}\right]\left[\frac{m_{\rm avg}}{0.3\ \text{\footnotesize$\rm M_{\odot}$}}\right]^{-1}\left[\frac{f_{\rm bin}}{0.5}\right]\times$ (7)
$\left[\frac{v_{\rm disp}}{100~\text{\footnotesize km/s}}\right]^{-1}\left[\frac{M_{\bullet}}{10^{7}~\text{\footnotesize$\rm M_{\odot}$}}\right]\left[\frac{r_{\rm close}}{200~\text{\footnotesize au}}\right]\ \text{yr}^{-1};$ (8)
in the lower expression, we have simplified the dependence on $f_{\rm bin}$ (formally valid for $f_{\rm bin}\ll 1$) and the gravitational focusing factor (since $f_{\text{grav}}\sim M_{\bullet}$ when $f_{\text{grav}}\gg 1$, which holds for all SMBHs and stars in their proximity). This value represents an upper bound, as the fiducial central density $\rho_{0}$ is near its maximum.
While the explicit dependence on parameters is instructive, and indeed is our choice in calculations of black hole growth (below), we also provide an expression with the full dependence on black hole mass, incorporating Equation (4), the $M_{\bullet}$-$v_{\rm disp}$ relation (Kormendy & Ho, 2013), and the potential variation of $r_{\rm close}$ with $M_{\bullet}$ (Hills, 1988):
$k\sim\frac{M_{\bullet}^{25/12}}{M_{b}^{2}+M_{\bullet}^{2}}\rightarrow\begin{cases}M_{\bullet}^{2.08}&\text{if }M_{\bullet}\ll M_{b},\\ M_{\bullet}^{0.08}&\text{if }M_{\bullet}\gg M_{b}.\end{cases}$ (9)
This equation suggests a “sweet spot” for binary break-up around $M_{\bullet}\approx M_{b}$, around $10^{7}$ $\rm M_{\odot}$. Next we explore how these results translate into black hole growth.

### v3

@@ Abstract
The growth of supermassive black holes (SMBHs) remains a central problem in astrophysics, with current observations providing only limited constraints on the underlying mechanisms. One possible growth channel is stellar accretion via the Hill’s mechanism, wherein a SMBH tidally breaks up a passing binary star, capturing and eventually accreting a member of the binary. We adopt a framework based on kinematics to predict capture rates from parameters that include the central number density of stars, the stellar velocity dispersion, the binary fraction, and black hole mass. We then estimate the growth of SMBHs across a range of galactic environments. In a data set of 91 galaxies of various types and masses, we identify two candidates with SMBHs for which stellar accretion may be a driver of growth. Closer to home, a recent analysis of observed hypervelocity stars from the Large Magellanic Cloud (LMC) implicates binary star interactions with a massive black hole. Every hypervelocity star produced in this way leaves a bound partner that may be accreted, providing an active growth channel for the LMC’s black hole.
@@ 1 Introduction
Enshrouding millions to billions of solar masses within an event horizon of remarkably small size, supermassive black holes are famously elusive. Yet observations have teased out irrefutable evidence for their existence. In our own galaxy, years-long tracking of stars on tight orbits about a common point beautifully reveal Sgr A, the Milky Way’s four million solar-mass black hole (Ghez et al., 2008; Genzel et al., 2010). The Event Horizon Telescope has since mapped gas flows near the supermassive black hole, unveiling a “shadow” surrounding the event horizon (Event Horizon Telescope Collaboration et al., 2022).
@@
Before these achievements were technologically feasible, J. Hills (Hills 1988) sought other circumstantial evidence for supermassive black holes. Focusing on the Milky Way, Hills calculated that a binary star passing close to a SMBH in the center of the Galaxy would be broken up, with one stellar component flung out at high speed in a gravitational sling-shot, while its partner would remain bound to the black hole. The unbound star would be traveling fast enough to escape the Galaxy. Detection of a hypervelocity star (HVS) produced by this binary disruption (hereafter, “break-up”) mechanism would be compelling evidence of the black hole itself.
@@
The discovery of a hypervelocity star by Brown et al. (2005) confirmed Hills’ prediction that a supermassive black hole in the center of the Milky Way produces fast stars with no other plausible origin. Subsequent detections (e.g., Brown et al., 2006, 2009; Zheng et al., 2014, see also Verberne et al. 2025) have built up a census of HVSs. The HVS population thus contributed to the growing body of evidence demonstrating that Sgr A was indeed a SMBH. More recently, based on Gaia astrometry (Gaia Collaboration et al., 2023), some of these same stars now appear to originate from the Large Magellanic Cloud (LMC). Han et al. (2025) analyzed these sources to infer the existence of LMC, a $\sim{6}\times 10^{5}$ $\rm M_{\odot}$ black hole in the dwarf galaxy. The Hills mechanism is fulfilling its promise as a signpost for SMBHs.
@@
Hypervelocity stars are the lead astronomical actors in the Hills mechanism, yet they are only half the story. The bound partner in each binary break-up remains close to the black hole and eventually may be accreted onto it (e.g., Antonini et al., 2010; Bromley et al., 2012; Bradnick et al., 2017; Yu & Lai, 2024; Sersante et al., 2025). A signature of this process is a tidal disruption event (TDE), which occurs when a black hole’s tidal forces shred the star (Hills, 1975; Frank & Rees, 1976; Lightman & Shapiro, 1977; Rees, 1988, see also Gezari 2021 and references therein). This outcome is possible for black holes with masses less than $\sim{10}^{8}$ $\rm M_{\odot}$, for which the typical tidal disruption radius is outside the event horizon. With more massive black holes, the star simply strays across the event horizon without a trace.
@@
Any bound star can be consumed, either tidally disrupted or swallowed by a black hole, once its orbital angular momentum drifts below a critical value, allowing it to plunge toward the event horizon (e.g., Merritt, 2013; Chen et al., 2023). As fuel for SMBHs, higher-mass stars are more susceptible to tidal disruption because of their diffuse outer layers, which experience weaker gravitational binding forces. As a result, when these stars approach an SMBH, their outer envelopes are easily stripped away by tidal forces (Chen et al., 2023). These partial disruptions can occur multiple times before a massive star is completely destroyed (Ryu et al., 2022). In any case, these stars, as well as their lower-mass counterparts, may ultimately be accreted by the SMBH (e.g., Thomsen et al., 2022; Webb & Foustoul, 2022; Olejak et al., 2025). Stellar accretion may indeed be a dominant early-stage growth channel for black holes in the mass range of $10^{2}$$10^{6}$ $\rm M_{\odot}$, including single-star captures (Milosavljević et al., 2006; Stone et al., 2017) and binary break-ups (González Prieto et al., 2025).
@@
A variety of mechanisms have been proposed to explain the growth of supermassive black holes, including mergers and gas accretion (e.g., Silk & Rees, 1998; Yu & Tremaine, 2002; Di Matteo et al., 2003; Montesinos Armijo & de Freitas Pacheco, 2011; Ryu et al., 2022). Yet no single channel has emerged to account for the evolution of all SMBHs. Here, our focus lies on the role of binary stars and the Hills mechanism as contributors to SMBH growth. We begin in §2 by revisiting the theoretical description of this growth channel from Bromley et al. (2012), implementing it in a Python code to track binary star interaction rates with a SMBH and the subsequent evolution of the black hole mass. Next, in §3, we apply our analysis to galactic data provided by Hannah et al. (2024, see also ), which span a wide range of black hole masses and galactic host properties. As a check on our methodology, we compare our approach to more sophisticated calculations of tidal disruption rates (§4). Closer to home, we assess the growth of LMC, the black hole candidate in the Large Magellanic Cloud S5. We summarize our results in §6.
@@ 2 The model: black hole growth through the Hills mechanism
The binary may be broken apart at larger closest-approach distances as well. Following Hills (1988, see also ), the closest-approach distance where disruptions are expected is
$r_{\rm close}$ $\approx$ $a_{\rm bin}\left(\frac{M_{\bullet}}{m_{\rm bin}}\right)^{1/3}$
$\approx$ $260\times\left[\frac{a_{\rm bin}}{1\ \text{\footnotesize au}}\right]\left[\frac{M_{\bullet}}{10^{7}~\text{\footnotesize$\rm M_{\odot}$}}\right]^{1/3}\left[\frac{m_{\rm bin}}{0.6\ \text{\footnotesize$\rm M_{\odot}$}}\right]^{-1/3}\text{au.}$
Hills (1988) noted that the break-up of a binary may occur at larger $r_{\rm close}$, depending on the binary’s orbital phase and orientation. Conversely, some binaries may survive the encounter with the SMBH intact at smaller $r_{\rm close}$, again, depending on the details of their orbits. [^1] Hills (1988) framed this range of outcomes probabilistically, with the likelihood of a break-up at $r_{\rm close}=0$ being 100% and decreasing linearly to zero at a closest-approach distance of $r_{{}_{\rm D}}/a_{\rm bin}$, where $r_{{}_{\rm D}}\approx r_{\rm close}(m_{\rm bin}/11M_{\bullet})^{1/3}$. Thus, when $r_{{}_{\rm D}}$ exceeds the binary separation (a value of about twice the distance in Eq. (2)) a binary can pass by a black hole unscathed. Nonetheless, Equation (2) provides a characteristic value of the binary break-up radius (Yu & Tremaine, 2003). Below, in calculating interaction cross sections, we adopt an even smaller fiducial value of $r_{\rm close}=200$ km/s, bringing the likelihood of a break-up closer to unity.
We also consider the conditions in which a binary break-up would yield a bound partner within the sphere of influence of the black hole,
$r_{v}=\frac{GM_{\bullet}}{v_{\rm disp}^{2}}\approx 4.3\times\left[\frac{M_{\bullet}}{10^{7}~\text{\footnotesize$\rm M_{\odot}$}}\right]\left[\frac{v_{\rm disp}}{100~\text{\footnotesize km/s}}\right]^{-2}\text{pc.}$ (2)
The semimajor axis of the bound partner following tidal break-up is
$a_{\rm bnd}$ $=$ $0.055\times\left[\frac{a_{\rm bin}}{1~\text{\footnotesize au}}\right]\left[\frac{m_{\text{ej}}}{0.3~\text{\footnotesize$\rm M_{\odot}$}}\right]^{-1}$
$\times\left[\frac{m_{\rm bin}}{0.6~\text{\footnotesize$\rm M_{\odot}$}}\right]^{1/3}\left[\frac{M_{\bullet}}{10^{7}~\text{\footnotesize$\rm M_{\odot}$}}\right]^{-1/3}\text{pc}$
where $m_{\text{ej}}$ is the mass of the ejected partner. This expression quantifies the limitations of binary disruption for wide binaries. In the case of a $10^{7}$ $\rm M_{\odot}$ SMBH, the bound partner of a binary with stars of mass 0.3 $\rm M_{\odot}$ would end up outside of the black hole’s sphere of influence if $a_{\rm bin}\gtrsim 80$ au. If a similar binary encountered a $10^{5}$ $\rm M_{\odot}$ black hole, neither star would be captured within the sphere of influence if the binary separation were much larger than 1 au.
@@ 2.1 Binary break-up rates
Here we are interested in representative binary properties to predict the change in mass of a SMBH from the accretion of stars captured by the Hills mechanism. Our starting point for the rate is $nv\sigma$” from kinematics, where $n$ is the number density of binaries, $v$ is their mean speed in a region surrounding a SMBH, and $\sigma$ is the cross-sectional area of the black hole — the size of the hypothetical target for a binary star to hit in order to be split into a hypervelocity star and a bound partner. We consider each factor in turn.
@@
where $m_{\rm avg}\approx$ 0.3 $\rm M_{\odot}$ is the average mass of a star, counting binary components as individual stars. In the Milky Way, the binary fraction is roughly half of all stars (Duchêne & Kraus, 2013, and references therein), though $f_{\rm bin}$ is dependent on the spectral type of the components (e.g., Moe & Di Stefano, 2017). Typically younger, more massive stars have a higher binary fraction than less massive, older stars.
The central mass density of stars, $\rho_{0}$, depends strongly on the SMBH mass, $M_{\bullet}$. For large elliptical galaxies, the density falls with increasing black hole mass $\rho_{0}\sim 1/M_{\bullet}$ (e.g., Faber et al., 1997). The trend is opposite in smaller galaxies. Both central stellar density and black hole mass increase with the size and mass of the galactic host. Here, we introduce the parameterization,
@@
where $\rho_{0}$ is the central density within some small (parsec-scale) distance of the SMBH, and $_{b}\approx 4\times 10^{7} M\rm M_{\odot}$ is the black hole mass associated with the break between two power-law relationships; $\rho_{0}\sim M_{\bullet}$ in the low-mass regime and $\rho_{0}\sim 1/M_{\bullet}$ for masses much greater than $M_{b}$.
Equation (5) is intended to provide representative values of the stellar mass density. Our parameterization in that equation was inspired by galaxy data kindly provided by Hannah et al. (2024), wherein the stellar mass density is evaluated at 5 pc away from the central SMBH. We demonstrate the effectiveness of Equation (5) in the next section. For a more detailed description of the stellar mass density in the centers of galaxies, we assume a power-law profile in radius $r$, so that
$\rho(r)=\rho_{0}\left(\frac{r}{r_{0}}\right)^{\gamma},$ (6)
where $\gamma\approx-2$ (Bahcall & Wolf, 1976) and we set $r_{0}=5$ pc. We make use of this expression in the next section as well.
The central velocity dispersion of stars in the cores of galaxies, $v_{\rm disp}$, is typically of the order of 100 km/s, with values that increase with SMBH mass and total galactic mass (e.g., Faber et al., 1997; Bernardi et al., 2003; McConnell & Ma, 2013; Greene et al., 2020). The range of observed values is narrow compared with that of the masses of known SMBHs, because the dispersion has weak dependence on black hole mass:
$v_{\rm disp}\propto M_{\bullet}^{1/4}$ (7)
(Kormendy & Ho, 2013) and broader galactic properties including luminosity (e.g., Faber & Jackson, 1976), which can span many orders of magnitude.
The third factor in the kinematic description is $\sigma_{\bullet}$, the cross-sectional area of the black hole in scattering events with binaries. This term depends on uncertain quantities including the distribution of binary orbital elements, and the nature of their orbits around the black hole. Following our review of the Hill mechanism above, we adopt $r_{\rm close}$ (Eq. (2)) for calculating the geometric cross-section near the black hole ($\pi r_{\rm close}^{2}$). Gravitational focusing by the SMBH boosts the geometric cross section by a factor of
@@
where the second term dominates for stellar orbits around supermassive black holes (Bondi & Hoyle, 1944). The aim of a binary toward the black hole may be wide, yet gravity draws it in.
Together, the three $nv\sigma$” terms provide an estimate of the rate of binary break-up through the Hills mechanism. In this kinematical approximation, the break-up rate of binary stars through the Hills mechanism is
$k$ $=$ $\frac{\rho_{0}}{m_{\rm avg}}\frac{f_{\rm bin}}{(1+f_{\rm bin})}\cdot v_{\rm disp}\cdot\pi f_{\text{grav}}r_{\rm close}^{2}$
$\approx$ $0.008\left[\frac{\rho_{0}}{10^{4}\ \text{\footnotesize$\rm M_{\odot}$/pc}^{3}}\right]\left[\frac{m_{\rm avg}}{0.3\ \text{\footnotesize$\rm M_{\odot}$}}\right]^{-1}\!\left[\frac{f_{\rm bin}}{0.1}\right]$
$\times\left[\frac{v_{\rm disp}}{100~\text{\footnotesize km/s}}\right]^{-1}\left[\frac{M_{\bullet}}{10^{7}~\text{\footnotesize$\rm M_{\odot}$}}\right]\left[\frac{r_{\rm close}}{200~\text{\footnotesize au}}\right]\ \text{yr}^{-1};$
in the lower expression, we have simplified the dependence on $f_{\rm bin}$ (formally valid for $f_{\rm bin}\ll 1$) and the gravitational focusing factor (since $f_{\text{grav}}\sim M_{\bullet}$ when $f_{\text{grav}}\gg 1$, which holds for all SMBHs and stars in their proximity). This value represents an upper bound, as the fiducial central density $\rho_{0}$ is near its maximum in Equation (5).
By considering the explicit SMBH mass dependence of the central density (Eq. (5)), the trend between black hole mass and velocity dispersion (Eq. (7)), and the interaction cross-section with gravitational focusing ($f_{\text{grav}}r_{\rm close}^{2}\propto M_{\bullet}^{5/6}$ (Hills, 1988), we estimate how the SMBH mass affects the binary break-up rate:
$k\propto\frac{M_{\bullet}^{25/12}}{M_{b}^{2}+M_{\bullet}^{2}}\rightarrow\begin{cases}M_{\bullet}^{2.08}&\text{if }M_{\bullet}\ll M_{b},\\ M_{\bullet}^{0.08}&\text{if }M_{\bullet}\gg M_{b}.\end{cases}$ (10)
This equation suggests a “sweet spot” of $M_{\bullet}\approx M_{b}$, around $10^{7}$ $\rm M_{\odot}$ where the fractional rate $k/M_{\bullet}$ reaches a peak. Next we explore how these results translate into black hole growth.